OpenStudy (anonymous):

Simplify. Express the product as a radical expression. ∛x² · fourth root of x The possible answers are a.7th root x^2 b.7th root x^3 c.12th root x^2 d.12th root x^11

OpenStudy (studygurl14):

\[\sqrt[3]{x^2}\times \sqrt[4]{x}\] Is that it?

OpenStudy (anonymous):

yes! @StudyGurl14

OpenStudy (jdoe0001):

\(\Large { \sqrt[3]{x^2}\cdot \sqrt[4]{x} \\ \quad \\ recall\implies \sqrt[{\color{red} m}]{a^{\color{blue} n}}=a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ \sqrt[3]{x^2}\cdot \sqrt[4]{x}\implies x^{\frac{2}{3}}\cdot x^{\frac{1}{4}}\implies \square ? }\)

OpenStudy (anonymous):

Also, just a little note, when you multiply exponents you're adding them.

OpenStudy (anonymous):

so is the answer b??? @iambatman

OpenStudy (jdoe0001):

is it? what did you get for \(\Large \sqrt[3]{x^2}\cdot \sqrt[4]{x}\implies x^{\frac{2}{3}}\cdot x^{\frac{1}{4}}\implies \square ?\)

OpenStudy (anonymous):

\[\huge x^{\frac{ 2 }{ 3 }+\frac{ 1 }{ 4 }}\]

OpenStudy (anonymous):

oh! nvm it would be 2/12... @jdoe0001

OpenStudy (anonymous):

so c

OpenStudy (jdoe0001):

\(\bf \cfrac{ 2 }{ 3 }+\cfrac{ 1 }{ 4 }\ne \cfrac{2}{12}\)

OpenStudy (anonymous):

@ImS0rryImN0t There seems to be a problem with fractions here for you, can you show us how you would add 2/3 + 1/4?